OpenStudy (anonymous):

how to substitute this x+y=16 x=7-5y

OpenStudy (anonymous):

really you just put your stuff for x in the equation and solve it you have to. since it is x + y = 16. the equations after you substitute it would be 7-5y+y = 16 then you can solve for y from there by combining like terms and blah blah

OpenStudy (anonymous):

can you show me the steps please

OpenStudy (anonymous):

x = 16-y Put that in the other equation (solve it in terms of y) 16-y = 7-5y 9 = -4y y = -9/4 Then put that back in the original equation x = 16-9/4 x = 55/4